OpenStudy (anonymous):

differentiate y=1/x anyone show working pliz how to solve

OpenStudy (anonymous):

With respect to x?

OpenStudy (gorv):

d/dx (x^n) = nx^(n-1) y = 1/x = x^(-1) dy/dx = -1 x^(-2) = - x^(-2) = -1/x^2 applying the power rule again, d^2 y/dx^2 = -1 * (-2 x^(-3)) = 2 x^(-3) = 2 /x^3 so u can calculate as mch as u want

OpenStudy (anonymous):

ok thanks